OpenStudy (anonymous):

(d^2y/dx^2)-5(dy/dx)+6y=xe^x

OpenStudy (anonymous):

i know that y_h(t)=c_1e^(3t)+c_2e^(2t) but now i have to find y_p(t) so to do this I said it equals Axe^x. therefore y'=A(x+1)e^x and y"=A(x+2)e^x and i plug those back into the original equation and try to solve for A but am getting stuck.

OpenStudy (anonymous):

This is a second order, linear, ordinary differential equation

OpenStudy (sidsiddhartha):

u can use P.I(proper integral) and C.F(complementary function) to solve this

OpenStudy (sidsiddhartha):

LHS can be written as\[(D^2-5D+6)y=0\] so the auxiliary eq is-- \[m^2-5m+6=0\\(m-3)(m-2)=0\\m=2,3\\CF=c_1e^{2x}+c_2e^{3x}\]

OpenStudy (sidsiddhartha):

now use this property-- \[\frac{ 1 }{ f(D) }e^{ax}V=e^{ax}\frac{ 1 }{ f(D+a) }V\]

OpenStudy (sidsiddhartha):

then proper integral will be simply-- \[PI=\frac{ 1 }{ D^2-5D+6 }(x*e^{x})\\=e^{x}*\frac{ 1 }{ (D+1)^2-5(D+1)+6 }x\] now try to simplyfy it

OpenStudy (unklerhaukus):

\[y''-5y'+6y=xe^x\] With \[y=Axe^x\]\[y' = A(1+x)e^x\]\[y'' = A(2+x)e^x\] Plugging these into the original DE\[A(2+x)e^x-5A(1+x)e^x+6Axe^x=xe^x\] Factoring this equation\[A\Big((2+x)-5(1+x)+6x\Big)e^x=(x)e^x\] \(e^x\) is not going to equal zero, so we can divide it out of the equation \[A\Big((2+x)-5(1+x)+6x\Big)=x\\ A\Big(2x-3\Big)=x\] Hmmm...

OpenStudy (unklerhaukus):

How did you get \(y_p(x) = Axe^x\)? try again with \(y_p=(Ax+B)e^x\) the derivatives are sure to be fun

OpenStudy (alekos):

I get A = 1/2 and B = 3/4 would anyone like to check?

OpenStudy (anonymous):

Thank you guys so much! It looks like I had the wrong set up for the y_p(t) equation! Haven't actually gone through the problem but in the answer key A=1/2 and B=3/4 like alekos said

OpenStudy (unklerhaukus):

thats it!